if a machine produce 1 car in 1 hr, how many cars would be produced by 100 machines in 100 hrs??
Answer by TimothyLamb(4379)

---
r = cars/(machine*hour) = 1/(1*1)
1/(1*1) = cars/(100 machine * 100 hour)
1 = cars/(100*100)
cars = 10000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php